OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 5i

OpenStudy (anonymous):

how to write the polynomial function in standard form with real coeffiecients whoses zeros include 1, 9i, -9i

OpenStudy (anonymous):

f(x)=(x-1)(x-9i)(x-(-9i))=(x-1)(x^2+81)=x^3-x^2+81x-81